OpenStudy (anonymous):

Simplify. n6 · n5 · n4 ÷ n3 · n2 ÷ n

OpenStudy (anonymous):

are there any parentheses?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

is n6 n^6 or 6n?

OpenStudy (anonymous):

step by step solution, needed?

OpenStudy (anonymous):

its n^6

OpenStudy (anonymous):

Cuz, it's pretty easy. in the exponents, there's this property that n^a. n^b= n^(a+b) So:- n^6 . n^5 .n^4 gives?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

well, i gave that property, right? a here is 6, b here is 5, and there's a c also, which is 4. so, if n^a . n^b . n^c = n^(a+b+c) that is???

OpenStudy (anonymous):

@triston69 go on, it's pretty easy, worst case scenario: you'll mess up. But you'll learn.

OpenStudy (anonymous):

There are no mark deductions on OS.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Dun' mention, but you figured out the solution. Right?

OpenStudy (anonymous):

ya